OpenStudy (zaarin):

the sum of 1/4 of the price of a pen and 1/3 of the price of a pencil is $11. If 1/8 of the price of the pen is equal to 1/5 of the price of the pencil, what is the price of the pencil in dollars? A.12 B.15 C.20 D.24 E.none of these

OpenStudy (anonymous):

system of equations: $11 = 1/4 * pen + 1/3 pencil 1/8 * pen = 1/5 * pencil

OpenStudy (zaarin):

I did that and I got yhe answer 6 is it correct?

OpenStudy (anonymous):

I can verify: eq1 + 2 x eq2 $11 = 1/4 * pen + 1/3 pencil +2/8 * pen = +2/5 * pencil $11 - 1/4 * pen= +5/15 pencil 1/4 * pen = 6/15 * pencil ------------------------ $11 = 11/15 * pencil $11 x 15 = 11 * pencil pencil = $15